A solid cube of sides 50 cm is kept in an empty reservoir of length 3 meters, breadth 2 meters and height 1 meter. The cube is taken out after filling the reservoir by water. What is the depth of water now?
Answer by mananth(16946)

Volume of cube = 0.5*0.5*0.5 =0.125 m^3
this volume is reduced after removing the cube
0.125 =3 *2*h
0.125/6 =h
it will reduce by 0.02m = 2 cm
1-0.02= 0.98 m height